40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8142     Entry time: Sat Feb 23 00:36:52 2013
Author: Manasa 
Type: Summary 
Category: Locking 
Subject: MC locked 

[Yuta, Manasa, Sendhil, Rana]

With MC REFL PD fixed, we aligned MC in high power enabling a fully functional MC autolocker.
We then unlocked MC and aligned the PD and WFS QPDs. Also we checked the MC demodulator and found a ~20% leakage between the Q-phase and I-phase. This must be fixed later by changing the cable length.

We adjusted MC offsets using /opt/rtcds/caltech/c1/scripts/MC/WFS/WFS_FilterBank_offsets.
We then measured the MC spot positions using  /opt/rtcds/caltech/c1/scripts/ASS/MC/mcassMCdecenter
Spot positions seem to have shifted by 2mm in yaw.

We will proceed with aligning the arms now.

Attachment 1: MCdecenter_23Feb2013.png  82 kB  | Hide | Hide all
MCdecenter_23Feb2013.png
ELOG V3.1.3-